A particular reactant decomposes with a half-life of 153 s when its initial concentration is 0.397 M. The same reactant decomposes with a half-life of 233 s when its initial concentration is 0.261 M Determine the reaction order. O 1 What is the value and unit of the rate constant for this reaction? Number s additional feedback
Answer #1

For n=2;
Units of K : M s-1
To find K use the formula below for two different values from the data,
ln (N2/N1)= K (t2-t1)
ln (200/100) = K (15 - 0)
K = 0.0462 min-1
Use this formula to find number of bacteria present after 135 min,
ln (N2 / 1600) = 0.0462 (135 - 60)
ln N2 - ln 1600 = 3.4657
ln N2 = 10.8435
N2 = e10.8435 = 51200 = bacteria present after 135 min